OpenStudy (anonymous):

factor the perfect square trinomial y^2-2/3y+1/9

Nnesha (nnesha):

fatima :) :) welcome to open study :0

OpenStudy (anonymous):

Thank u :)

OpenStudy (aum):

To complete the square for x^2 + bx + c, you divide the coefficient of the middle term by 2 to get b/2 and then complete the square as (x - b/2)^2. y^2-2/3y+1/9 Coefficient of the middle term is -2/3. Divide it by 2: -2/3 divided by 2 = -1/3. (y - 1/3)^2 will complete the square. (y - 1/3)^2 = y^2 - 2/3y + 1/9

OpenStudy (aum):

Here they tell you it is a perfect square and so we can stop after doing the above. If it is not a perfect square you have to include what remains after completing the square.